OpenStudy (anonymous):

dy/dx=1+x solve by differential equation

OpenStudy (amistre64):

y' = 1+x y = x + (1/2)x^2 + C Does that fit?

OpenStudy (anonymous):

\[y(x) = c_1+1/2 x (x+2)\]

OpenStudy (amistre64):

y = x[(x/2)+1] + C ??

OpenStudy (anonymous):

\[dy = \left( x + 1 \right)dx\] \[\int\limits_{}{dy}=\int\limits_{}(x+1){dx}\] \[y= (1/2) x^2 + x + C\]